# Optimizing energy conversion with nonthermal resources in steady-state quantum devices

### v1

@@ IV Optimized cooling exploiting a nonthermal resource
Figure 7: Optimized performance for concrete nonthermal distributions. (a) Mixed distribution, Eq. (44) and its thermal equivalent, both used to obtain (b) optimal efficiency (c) optimally low noise-to-signal ratio and (d) precision-efficiency trade-off. (e) Zoom-in of (a) where non-cooling regions of the nonthermal distribution are shaded in gray. (f) to (g) show optimal transmission functions at the fixed cooling power indicated by crosses and stars in (b) to (d). (i) Dip-peak distribution, Eq. (45) and its thermal equivalent, both used to obtain (j) optimal efficiency (k) optimally low noise-to-signal ratio and (l) precision-efficiency trade-off. (m) Zoom-in of (i) where non-cooling regions of the nonthermal distribution are shaded in gray. (n) to (p) show optimal transmission functions at the fixed cooling power indicated by crosses and stars in (j) to (l).
We have until here introduced an optimization procedure for general output currents, resources, efficiencies and precision for a generic steady-state two-terminal coherent conductor described by scattering theory. In this section, we apply the developed procedure to a relevant thermodynamic process, namely cooling. The task to be completed is hence to achieve a positive steady-state cooling power $I^{Q,\mathrm{L}}$ such that $I^{x}\to I^{Q,\mathrm{L}}/T$, while using the nonthermal distribution in the right contact as resource and the transmission properties of the quantum conductor as energy filter. The resource “used up” can be quantified by the entropy production in the right (nonthermal) reservoir $I^{y}\to-I^{\Sigma,\mathrm{R}}$. The sketch of Fig. 1 illustrates the working principle.
@@
In order to compare the opportunities arising from a nonthermal distribution as a resource to those from a thermal resource, we use the concept of Büttiker probes [57, 58, 59, 60, 61] to assign an effective temperature $T_{\mathrm{eff}}$ and an effective electrochemical potential $\mu_{\mathrm{eff}}$ to the probe distribution. A Büttiker probe can be imagined as a floating contact attached to the terminal with the nonthermal distribution, where the probe adjusts its temperature and electrochemical potential under the condition that on average no particles and no energy is transferred between the nonthermal terminal and the probe  [^2] We choose a standard measure to define the effective temperature and chemical potential to the average charge and energy current while other measures might consider, e.g., the noise in the particle current. See [8] for an overview on thermometry on the nanoscale..
In the following, we analyze the performance of cooling with the two example nonthermal distributions. We determine the series of boxcar transmissions that is required to optimize the performance as quantified by cooling power, efficiency, precision and a trade-off relation. We finally show in Sec. IV.3 how the performance is modified, when the ideal boxcar transmissions are replaced by Lorentzian transmissions, which is one commonly occurring transmission function stemming from resonances in experimentally realized conductors. For details of the numerical procedures employed, see Appendix D and Ref. [62].
@@ IV.1 Nonthermal distribution from competing environments
We start by presenting an example of a nonthermal distribution consisting of a mixture of two thermal distributions, henceforth referred to as a “mixed distribution”. Such a distribution can result from the competition of different environments (such as sun light and phonon background in solar cells [19] or from two mixed electronic distributions at different potentials in a quantum Hall setup [26, 24], to name two examples). We construct the nonthermal distribution from
@@
with $a+b=1$ and with the thermal distributions of the competing environments $f_{\mathrm{env},i}(\varepsilon)$ at temperatures $T_{\mathrm{env},i}$ and electrochemical potentials $\mu_{\mathrm{env},i}$. The resulting distribution, $g(\varepsilon)$, is shown as purple line in Fig. 7(a) for $a=b=1/2$ and for $\mu_{\mathrm{env},1}=-2k_{\mathrm{B}}T$ and $T_{\mathrm{env},1}=5T$, $\mu_{\mathrm{env},2}=-0.5k_{\mathrm{B}}T$ and $T_{\mathrm{env},2}=1.2T$ , where the nonthermal nature is seen in the “step” which is not present in thermal distributions. The hot thermal distribution that is found from connecting the nonthermal reservoir to a Büttiker probe has $T_{\mathrm{eff}}=3.5035\>T$ and $\mu_{\mathrm{eff}}=-1.2214\>k_{\mathrm{B}}T$ (rounded to the fourth decimal), and it is shown as the red line in the same panel. The cold distribution in the left contact has temperature $T$ and electrochemical potential $\mu=0$.
@@ IV.2 Nonthermal distribution from irradiation
We furthermore consider an example distribution consisting of an underlying thermal distribution with an added Lorentzian peak and dip, see Fig. 7(i). Such a dip-peak shape mimics excitations stemming from fixed-frequency light irradiation, promoting electron excitations to higher energies while leaving behind hole-like excitations in the Fermi sea, see for example Refs. [63, 64]. This distribution has the interesting feature of nonmonotonicity and it will be referred to as the “dip-peak distribution” henceforth. Concretely, it is defined by
@@ IV.2.1 Optimization by tuning the transmission function
We start by showing the results for the optimal efficiency, noise-to-signal ratio, and trade-off parameter between efficiency and precision at fixed cooling power in Fig. 7(j)-(l) obtained by optimizing the transmission function $\mathcal{D}(\varepsilon)$ in analogy to the study based on the mixed distribution in Sec. IV.1. Similar to the results achieved with the mixed distribution, the dip-peak distribution consistently performs better than its thermal equivalent for each of the quantifiers. In particular the improvement concerning the precision and trade-off in panels (k) and (l) sticks out.
This can again be understood by analyzing the transmission functions that optimize the performance as shown in panels (n) to (p), for the fixed current $I^{\Sigma,\mathrm{L}}=0.0008\>k_{\mathrm{B}}^{2}T/h$. A crucial difference compared to the results obtained with the mixed distribution results from the nonmonotonicity of the dip-peak distribution: the optimal transmission functions therefore now consist of series of boxcar functions—instead of only a single one.
@@
The additional crossing points showcase that nonthermal distributions can give additional freedom to the optimization. The optimal transmission functions for the thermal distribution are practically defined by fewer variables: the optimal function for efficiency always starts at the crossing point while for precision it always starts the at the electrochemical potential, the farthest possible points from the crossing points. Thus, only the other end of the boxcar is determined by the fixed output current. However, for the nonthermal distribution, there are two additional crossing points. The optimal transmission function in panels (n) and (m) picks out areas near these, while keeping the transmission window narrow in the cooling window the nonthermal distribution and its thermal equivalent share. The trade-off optimization also exploit the new cooling windows around the nonmonotonicity.
@@ IV.2.2 Optimization with $\mu$ as second variable
Figure 8: (a) Optimal efficiency, obtained from the nonthermal dip-peak distribution and from its thermal equivalent by optimizing the transmission function $\mathcal{D}(\varepsilon)$ as well as the electrochemical potential of the cold thermal distribution $f(\varepsilon)$, as function of the fixed cooling power $I^{\Sigma,\mathrm{L}}$. (b) Optimal efficiency as function of $\mu$, which optimizes the efficiency at each fixed cooling power. (c) Relation between fixed cooling power $I^{\Sigma,\mathrm{L}}$ and the electrochemical potential $\mu$ leading to optimal efficiency at $I^{\Sigma,\mathrm{L}}$.
@@ IV.3 Approaching boxcar transmission with Lorentzians
We have until here found ideal realizations of the transmission functions—namely series of sharp boxcar-shaped transmissions—that optimize the performance of the cooling. These results are important as they set bounds on how good the cooling power, the efficiency at fixed cooling power, or the precision at fixed cooling power can get. Realistically, the transmission function in a quantum system is however not expected to be of ideal boxcar shape—fine tuning of for example a large amount of quantum dots coupled in series or similar band engineering would be required [33, 11]. The question arises hence how the performance will be impacted and how close to the bounds one gets, when the transmission function of the conductor only approaches this ideal boxcar shape.
Figure 10: (a) Efficiency at finite output power for a series of optimized boxcars and approximations by Lorentzians as shown in panel (b). Results for the nonthermal distribution as well as its thermal equivalent are shown in purple and red. (b) Transmission function consisting of three boxcars optimising efficiency at fixed output power for the distribution function of Eq. (45) (light purple) and series of Lorentzians approximating the ideal case (dark purple).
Concretely, we here show how the performance will be influenced when the boxcar-shaped transmissions are replaced by Lorentzian-shaped transmissions
@@
with maximum transmission $\mathcal{D}_{\mathrm{Lor},j}$ found at $\varepsilon=E_{j}$ and with half width at half maximum $\Gamma_{j}$. Such Lorentzian-shaped transmissions are realized in many experimental settings where resonances occur [65], such as in low-temperature experiments with gate-tunable quantum dots [66], in molecular electronics even at room temperature [67, 68], or in electronic Fabry-Perot interferometers [69, 70]. The full transmission of this realistic example then has the form of a series of resonances
@@
We here choose to adapt this series of resonances to the ideal series of boxcar-shaped transmission windows by placing the center of the resonances at the center of the boxcars, constituting the optimal transmission $\mathcal{D}_{\mathrm{opt}}(\varepsilon)$. [^4] This should not be confused with an optimization of the efficiency under the constraint that the transmission is a series of Lorentzians, which is not the goal here. Although, the results of such an optimization would probably be close to the chosen model. Then, we choose to fix the widths of the Lorentzian transmissions to $0.1\times$ the width of a given boxcar-shaped part of the optimal transmission. An example for this is shown in Fig. 10(b). In this way most of the Lorentzian is contained inside the boxcar transmission that would constitute the ideal case.
@@
Finally, note that there is a large variety of energy-dependent transmission functions in different nanostructures, where we here only analyze one example model. By only using one Lorentzian per boxcar of the ideal transmission, the output power is severely lowered due to the necessity of narrow peaks. However, for more complex systems, like a larger and more accurate combination of quantum dots, the realistic transmission function is expected to approach a boxcar shape more closely. What our cursory investigation conveys is that the optimal transmission function can be used to guide the design of realistic functions: the closer they are to the optimal boxcars, the closer the performance should be to the bound.
@@ V Conclusion
We have developed a general approach to optimize the performance of energy conversion processes in two-terminal coherent conductors operating in steady-state, focusing on devices fed by nonthermal distributions and that can be modeled by scattering theory. We therefore extended an approach based on Lagrange multipliers put forward in Refs. [33, 11] to generic performance quantifiers—such as general efficiencies, precision, and trade-off relations—and to the nonthermal case.
@@ Appendix A Crossing point illustrations
As a next step, consider $\lambda$ in the condition functions. Once the expressions for the condition functions, and thus the respective optimal transmission functions, are found, we can consider the output current as a function of $\lambda$, $I^{x}=I^{x}(\lambda)$. By regarding $\lambda$ as a variable, the spectrum of possible $I^{x}$ is found by evaluating the current integral with the optimal transmission functions. The minimum and maximum currents are found in the limits of $\lambda\in[0,\infty)$. As $\lambda\to\infty$, the output current reaches the maximum possible current. This means that the optimal transmission function reaches the transmission function for maximum output current. Furthermore, the condition functions for the optimized quantities must converge to the one for maximum cooling power. How close the functions are to the maximum condition function depends on $\lambda\in[0,\infty)$, where larger values means that the optimal transmission functions encompass larger areas of the cooling window(s).
@@ Appendix B Noise minimization based on thermal-like contribution only
$f(\alpha x+(1-\alpha)y)\geq\alpha f(x)+(1-\alpha)f(x),\alpha\in[0,1]$ (48)
@@
where we introduced the constants $a_{\gamma}$ and $b_{\gamma}$ for a certain slice $\gamma$ comparing to the prefactors of $d_{\gamma}$ and $d_{\gamma}(1-d_{\gamma})$ in Eq. (B). This function is clearly a concave hyperbolic function, due to the negative square term. When integrating, one sums over all the slices $\gamma$ and takes the limit $\delta\epsilon\rightarrow 0$. Since the sum is a concave function, we can conclude that the integral $S^{x}$ is a concave functional of $\mathcal{D}(\varepsilon)$ which implies that there is no local minimum and the global minimum is on the bounds of the function domain, meaning that the transmission must be either zero or one at every energy for the global minimum solution. A more rigorous argument for this can be found in [34].
@@ Appendix C Classification of spectral currents
Only two of the four regimes presented in Table 1 are in principle supposed to be relevant for the optimization: the regular and the reversed operation. The other two are either forbidden or not beneficial. The forbidden region refers to a set of conditions under which it should be impossible for both a resource and an output to be generated simultaneously, assuming the currents are well-defined and comply with the second law of thermodynamics. If such a region were to exist, the transmission function could be defined exclusively within it, leading to an unphysical system. For instance, consider a case where the resource current corresponds to the entropy production in one reservoir and takes a negative value, while the output current represents the entropy reduction in another reservoir (i.e., negative entropy production). The presence of a forbidden region would then imply the possibility of constructing a device with negative total entropy production, as entropy would be reduced in both reservoirs.
The non-beneficial region, on the other hand, can exist. Using the previous example, the entropy would increase in both reservoirs, which is physically allowed. However, this regime provides no possible benefit to the performance and is hence excluded from a device operation that is supposed to have optimal efficiency.

### v2

@@ IV Optimized cooling exploiting a nonthermal resource
Figure 7: Optimized performance for concrete nonthermal distributions. (a) Mixed distribution, Eq. (45) and its thermal equivalent, both used to obtain (b) optimal efficiency (c) optimally low noise-to-signal ratio and (d) precision-efficiency trade-off. (e) Zoom-in of (a) where non-cooling regions of the nonthermal distribution are shaded in gray. (f) to (h) show optimal transmission functions at the fixed cooling power indicated by crosses and stars in (b) to (d). (i) Dip-peak distribution, Eq. (46) and its thermal equivalent, both used to obtain (j) optimal efficiency (k) optimally low noise-to-signal ratio and (l) precision-efficiency trade-off. (m) Zoom-in of (i) where non-cooling regions of the nonthermal distribution are shaded in gray. (n) to (p) show optimal transmission functions at the fixed cooling power indicated by crosses and stars in (j) to (l).
We have until here introduced an optimization procedure for general output currents, resources, efficiencies and precision for a generic steady-state two-terminal coherent conductor described by scattering theory. In this section, we apply the developed procedure to a relevant thermodynamic process, namely cooling. The task to be completed is hence to achieve a positive steady-state cooling power $I^{Q,\mathrm{L}}$ such that $I^{x}\to I^{Q,\mathrm{L}}/T$, while using the nonthermal distribution in the right contact as resource and the transmission properties of the quantum conductor as energy filter. The resource “used up” can be quantified by the entropy production in the right (nonthermal) contact $I^{y}\to-I^{\Sigma,\mathrm{R}}$. The sketch of Fig. 1 illustrates the working principle.
@@
In order to compare the opportunities arising from a nonthermal distribution as a resource to those from a thermal resource, we use the concept of Büttiker probes [Buttiker1985Aug, Buttiker1986Mar, Buttiker1988Nov, deJong1996Aug, Forster2007Jan] to assign an effective temperature $T_{\mathrm{eff}}$ and an effective electrochemical potential $\mu_{\mathrm{eff}}$ to the probe distribution. A Büttiker probe can be imagined as a floating contact attached to the terminal with the nonthermal distribution, where the probe adjusts its temperature and electrochemical potential under the condition that on average no particles and no energy is transferred between the nonthermal terminal and the probe  [^2] We choose a standard measure to define the effective temperature and chemical potential to the average charge and energy current while other measures might consider, e.g., the noise in the particle current. See [Giazotto2006Mar] for an overview on thermometry on the nanoscale..
In the following, we analyze the performance of cooling with the two example nonthermal distributions. We determine the series of boxcar transmissions that is required to optimize the performance as quantified by cooling power, efficiency, precision and a trade-off relation. We finally show in Sec. IV.3 how the performance is modified, when the ideal boxcar transmissions are replaced by Lorentzian transmissions, which is one commonly occurring transmission function stemming from resonances in experimentally realized conductors. For details of the numerical procedures employed, see Appendix D and Ref. [zenodo].
@@ IV.1 Nonthermal distribution from competing environments
We start by presenting an example of a nonthermal distribution consisting of a mixture of two thermal distributions, henceforth referred to as a “mixed distribution”. Such a distribution can result from the competition of different environments (such as sun light and phonon background in solar cells [Tesser2023] or from two mixed electronic distributions at different potentials in a quantum Hall setup [leSueur2010Jul, Altimiras2010Nov], to name two examples). We construct the nonthermal distribution from
@@
with $a+b=1$ and with the thermal distributions of the competing environments $f_{\mathrm{env},i}(\varepsilon)$ at temperatures $T_{\mathrm{env},i}$ and electrochemical potentials $\mu_{\mathrm{env},i}$. The resulting distribution, $g(\varepsilon)$, is shown as purple line in Fig. 7(a) for $a=b=1/2$ and for $\mu_{\mathrm{env},1}=-2k_{\mathrm{B}}T$ and $T_{\mathrm{env},1}=5T$, $\mu_{\mathrm{env},2}=-0.5k_{\mathrm{B}}T$ and $T_{\mathrm{env},2}=1.2T$ , where the nonthermal nature is seen in the “step” which is not present in thermal distributions. The hot thermal distribution that is found from connecting the nonthermal contact to a Büttiker probe has $T_{\mathrm{eff}}=3.5035\>T$ and $\mu_{\mathrm{eff}}=-1.2214\>k_{\mathrm{B}}T$ (rounded to the fourth decimal), and it is shown as the red line in the same panel. The cold distribution in the left contact has temperature $T$ and electrochemical potential $\mu=0$.
@@ IV.2 Nonthermal distribution from irradiation
We furthermore consider an example distribution consisting of an underlying thermal distribution with an added Lorentzian peak and dip, see Fig. 7(i). Such a dip-peak shape mimics excitations stemming from fixed-frequency light irradiation, promoting electron excitations to higher energies while leaving behind hole-like excitations in the Fermi sea, see for example Refs. [Song2015Apr, Kim2011Aug]. This distribution has the interesting feature of nonmonotonicity and it will be referred to as the “dip-peak distribution” henceforth. Concretely, it is defined by
@@ IV.2.1 Optimization by tuning the transmission function
We start by showing the results for the optimal efficiency, noise-to-signal ratio, and trade-off parameter between efficiency and precision at fixed cooling power in Fig. 7(j)-(l) obtained by optimizing the transmission function $\mathcal{D}(\varepsilon)$ in analogy to the study based on the mixed distribution in Sec. IV.1. Similar to the results achieved with the mixed distribution, the dip-peak distribution consistently performs better than its thermal equivalent for each of the quantifiers. In particular the improvements concerning the precision and trade-off in panels (k) and (l) stick out.
This can again be understood by analyzing the transmission functions that optimize the performance as shown in panels (n) to (p), for the fixed current $I^{\Sigma,\mathrm{L}}=0.0008\>k_{\mathrm{B}}^{2}T/h$. A crucial difference compared to the results obtained with the mixed distribution results from the nonmonotonicity of the dip-peak distribution: the optimal transmission functions now consist of series of boxcar functions—instead of only a single one.
@@
The additional crossing points showcase that nonthermal distributions can give additional freedom to the optimization. The optimal transmission functions for the thermal distribution are practically defined by fewer variables: the optimal function for efficiency always starts at the crossing point while for precision it always starts the at the electrochemical potential, the farthest possible point from the crossing point. Thus, only the other end of the boxcar is determined by the fixed output current. However, for the nonthermal distribution, there are two additional crossing points. The optimal transmission function in panels (n) and (m) picks out areas near these, while keeping the transmission window narrow in the cooling window shared by the nonthermal distribution and its thermal equivalent. The trade-off optimization also exploit the new cooling windows around the nonmonotonicity.
@@ IV.2.2 Optimization with $\mu$ as second variable
[…]
@@ IV.3 Approaching boxcar transmission with Lorentzians
We have until here found ideal realizations of the transmission functions—namely series of sharp boxcar-shaped transmissions—that optimize the performance of the cooling. These results are important as they set bounds on how good the cooling power, the efficiency at fixed cooling power, or the precision at fixed cooling power can get. Realistically, the transmission function in a quantum system is however not expected to be of ideal boxcar shape—fine tuning of for example a large amount of quantum dots coupled in series or similar band engineering would be required [Whitney2014Apr, Whitney2015]. The question arises hence how the performance will be impacted if the transmission differs from this ideal case. To achieve boxcar-like transmissions one could for example employ quantum dot chains [Cha2020Oct, Ehrlich2021] together with dedicated tuning of the dot properties [Kalantre2019Jan], or exploit metal-molecule-metal junctions where exchanging anchoring groups of the central molecule shapes the electronic transmission [Tsukamoto2012].
Here, instead of analysing specific, complex realizations potentially approaching a boxcar transmission, we address the example of a transmission function that is also limited in energy but standardly achieved in experiments. Namely, we analyze a Lorentzian transmission
Figure 10: (a) Efficiency at finite output power for transmission functions consisting of a series of Lorentzians as exemplified in panel (b). Results for the nonthermal distribution as well as its thermal equivalent are shown in purple and red. (b) Transmission function consisting of three boxcars optimizing efficiency at a fixed output power, indicated by the purple plus in panel (a), for the distribution function of Eq. (46) (light purple) and series of Lorentzians approximating the ideal case (dark purple).
@@
with maximum transmission $\mathcal{D}_{\mathrm{Lor},j}$ found at $\varepsilon=E_{j}$ and with half width at half maximum $\Gamma_{j}$. Indeed, such Lorentzian-shaped transmissions are realized in many experimental settings where resonances occur [Ihn2009Dec], such as in low-temperature experiments with gate-tunable quantum dots [Kouwenhoven2001Jun], in molecular electronics even at room temperature [Xin2019Mar, Gemma2023Jun], or in electronic Fabry-Perot interferometers [Deprez2021May, Camino2007Oct]. The full transmission of this realistic example has the form of a series of resonances
@@
We here choose to adapt this series of resonances to the ideal series of boxcar-shaped transmission windows by placing the center of the resonances at the center of the boxcars, constituting the optimal transmission $\mathcal{D}_{\mathrm{opt}}(\varepsilon)$. [^4] This should not be confused with an optimization of the efficiency under the constraint that the transmission is a series of Lorentzians, which is not the goal here. Then, we choose to fix the widths of the Lorentzian transmissions to $0.1\times$ the width of a given boxcar-shaped part of the optimal transmission. An example for this is shown in Fig. 10(b). In this way most of the Lorentzian is contained inside the boxcar transmission that would constitute the ideal case.
@@
Finally, note that there is a large variety of energy-dependent transmission functions in different nanostructures, while we here only analyze one example model. By only using one Lorentzian per boxcar of the ideal transmission, the output power is severely lowered due to the necessity of narrow peaks. However, for more complex systems, like a larger and more accurate combination of quantum dots, the realistic transmission function is expected to approach a boxcar shape more closely. What our cursory investigation conveys is that the optimal transmission function can be used to guide the design of realistic functions: the closer they are to the optimal boxcars, the closer the performance should be to the bound.
@@ V Conclusion
We have developed a general approach to optimize the performance of energy conversion processes in two-terminal coherent conductors operating in steady-state, focusing on devices fed by nonthermal distributions and that can be modeled by scattering theory. We therefore extended an approach based on Lagrange multipliers put forward in Refs. [Whitney2014Apr, Whitney2015] to generic performance quantifiers—such as general efficiencies, precision, and trade-off relations—and to the nonthermal case.
@@ Acknowledgements.
Data availability
The data that support the findings are openly available [zenodo].
@@ Appendix A Crossing point illustrations
As a next step, consider $\lambda$ in the condition functions. Once the expressions for the condition functions, and thus the respective optimal transmission functions, are found, we can consider the output current as a function of $\lambda$, $I^{x}=I^{x}(\lambda)$. By regarding $\lambda$ as a variable, the spectrum of possible $I^{x}$ is found by evaluating the current integral with the optimal transmission functions. The minimum and maximum currents are found in the limits of $\lambda\in[0,\infty)$. As $\lambda\to\infty$, the output current reaches the maximum possible current. This means that the optimal transmission function reaches the transmission function for maximum output current. Furthermore, the condition functions for the optimized quantities must converge to the one for maximum cooling power. How close the functions are to the maximum condition function depends on $\lambda\in[0,\infty)$, where larger values mean that the optimal transmission functions encompass larger areas of the cooling window(s).
@@ Appendix B Noise minimization based on thermal-like contribution only
$f(\alpha x+(1-\alpha)y)\geq\alpha f(x)+(1-\alpha)f(y),\alpha\in[0,1]$ (49)
@@
where we introduced the constants $a_{\gamma}$ and $b_{\gamma}$ for a certain slice $\gamma$ comparing to the prefactors of $d_{\gamma}$ and $d_{\gamma}(1-d_{\gamma})$ in Eq. (B). This function is clearly a concave hyperbolic function, due to the negative square term. When integrating, one sums over all the slices $\gamma$ and takes the limit $\delta\epsilon\rightarrow 0$. Since the sum is a concave function, we can conclude that the integral $S^{x}$ is a concave functional of $\mathcal{D}(\varepsilon)$ which implies that there is no local minimum and the global minimum is on the bounds of the function domain, meaning that the transmission must be either zero or one at every energy for the global minimum solution. A more rigorous argument for this can be found in [Landi2025].
@@ Appendix C Classification of spectral currents
Only two of the four regimes presented in Table 1 are in principle supposed to be relevant for the optimization: the regular and the reversed operation. The other two are either forbidden or not beneficial. The forbidden region refers to a a set of conditions under which both a resource and an output are generated simultaneously, which should be impossible assuming the currents are well-defined and comply with the second law of thermodynamics. If such a region were to exist, the transmission function could be defined exclusively within it, leading to an unphysical system. For instance, consider a case where the resource current corresponds to the entropy production in one contact and takes a negative value, while the output current represents the entropy reduction in another contact (i.e., negative entropy production). The presence of a forbidden region would then imply the possibility of constructing a device with negative total entropy production, as entropy would be reduced in both contacts.
The non-beneficial region, on the other hand, can exist. Using the previous example, the entropy would increase in both contacts, which is physically allowed. However, this regime provides no possible benefit to the performance and is hence excluded from a device operation that is supposed to have optimal efficiency.